Bank Information

Rhiza S. Sadjad

Control Systems and Instrumentation Laboratory, Department of Electrical Engineering Faculty of Engineering, Hasanuddin University, MAKASSAR Indonesia 90245 Email: rhiza@unhas.ac.id, URL: http://www.unhas.ac.id/rhiza/

I. INFORMATION

Bank Name: MIZUHO BANK, LTD O-okayama Branch Address: 8-28-12 KITASENZOKU OHTA-KU TOKYO JAPAN 145-0062 Swift BIC: MHBKJPJT Account No. : 145-2242911 Name: SADJAD RHIZA SJAMSOEOED Address: International House # 222 1-1-18 Ishikawa-Cho OTA-KU TOKYO JAPAN 145-0061 Mobile: +81 90 6474 0609

II. COPY of the SAVING BOOK

Fig. 1. The Copy of the Saving Account Book